### v1

@@ 4 Results and Analysis
In this section, we present the results of our best training setup over the full field of the selected SDC1 image. We stress that there are small variations of results from retraining a network using the same training setup, with typical variations of the best achievable score around ±$\pm 0.5\%$. In practice, the detailed analyses carried out on this specific training hold perfectly with any training using the same setup. The results in this section are ordered following the post-processing pipeline from the raw predicted quantities up to the predicted source catalog using the challenge format. The validation loss curve as a function of the training iteration is presented in Fig. 13. This figure also indicates the optimized score at a regular interval of 200 iterations. The following results are presented for iteration 3600, which achieved the best score for this specific training.
@@ 4.1 Patch-divided prediction results
From the raw model prediction, after applying the first objectness filtering and the in-patch NMS, we can obtain a list of all the detected sources per patch. We represent the respective number of sources for all patches over the full image field in Fig. 14. We note that some sources are counted multiple times in this representation because the inter-patch NMS has not yet been applied. We see that the detector sensitivity follows the primary beam imprints over the full image. We do not observe a clear overdensity of detected sources in the training area region, which would be a clear sign of overtraining.
To analyze in detail the behavior of the individual detection units from our detector output (Sect. 2.5), we represent a histogram of the number of found-useful detections per grid cell for all patches in Fig. 15. Here, empty grid cells are excluded, but they represent the more common case, especially considering that we keep patches far from the image center. We observe that cases with 6 or more detections are very few, 4 or 5 detections are not common but significant, and most non-empty grid cells contain less than 3 detections. Further analysis indicates that only around $26\%$ of the non-empty grid cells have more than 1 detection, and this proportion drops to $4.5\%$ for more than 2 detections. In fact, the in-patch NMS process removes $78\%$ of the raw good detections. This indicates that many detection units are in use and capable of detecting real sources, but they often end up detecting the same sources. Still, we observed that lowering the number of detection units lowers the score, indicating that they are useful for some specific contexts or that having more detection units better distributes the information. To further verify this idea, we represent in Fig. 16 a histogram of the number of predictions per detection unit for the whole image. The distribution over the different detection units from the small-size regime is relatively homogeneous, confirming that all detection units are useful. We also observe that the relative proportions of each size regime are well preserved regarding the theoretical target distribution from Fig. 12, which confirms that our association process properly distributes the information over the available size regimes.
To have a better view of the representation regime of individual detection units, we can look at their respective distribution of predicted sizes and aspect ratios in Fig. 17. We observe that some boxes are more specialized regarding aspect ratios, which is natural for the intermediate size regime due to the two different ratios of size priors. The behavior of the small-size regime detection units is more surprising, as it appears to prefer a positive aspect ratio for small but resolved sources. The cause of this behavior remains unidentified, but we can firmly state that it does not come from the training sample for which the target box aspect ratios are well balanced. Finally, the detection unit from the large-size regime and usually one or two from the small-size regime predict boxes with actual sizes that would be more representative of the intermediate regime. Due to the significant difference in aspect ratio, only one of the intermediate-size regime detection units is likely to be used in a given grid cell. In cases where two or more sources of intermediate size have to be detected in the same grid cell, a small or large box with a square prior might be preferable to using the wrong aspect ratio box of the appropriate size regime. This effect is less visible if we train our detector using only square priors.
Figure 15: Distribution of the non-empty grid cells regarding the number of detections in each cell, after objectness filtering and in-patch NMS.
@@
Figure 17: Histograms of the predicted heights, widths, and aspect ratios for every detection unit. The aspect ratio is computed as $(w-h)/\max{(w,h)}$. The total number of predictions and the proportion of strictly positive and negative aspect ratios are also given for each detection unit. The red vertical lines indicate the width and height of the size priors.
@@ 4.2 Full catalog scoring
Table 5: . List of SDC1 scores and related properties for source catalogs from different teams and methods.
Team (method) $M_{s}$ (Score) $N_{det}$ $N_{match}$ $N_{false}$ $N_{bad}\in N_{false}$ Purity $\bar{s}$
@@
MINERVA (YOLO-CIANNA-ref) 479758 718760 677025 41735 15787 94.19% 0.7703
MINERVA (YOLO-CIANNA-alt) 414937 538649 533659 4990 2402 99.07% 0.7869
@@
YOLO-CIANNA-ref
YOLO-CIANNA-alt
Figure 18: Histograms of the sources as a function of their flux using logarithmic bins for the underlying True catalog and the predicted sources in the testing area. The true source flux is used for matches, while the predicted flux is used for false positives. The bottom part of each frame represents the purity and completeness of each bin of the histograms (defined in Sect. 4.2). The top frames use the apparent flux after multiplication with the primary beam, while the bottom frames use the integrated flux. The left column represents the results for the YOLO-CIANNA-ref model, and the right column the YOLO-CIANNA-alt model.
After the inter-patch NMS filtering and conversion of all the predicted source parameters, we obtain a source catalog for the full image that can be put through the SDC1 scorer (Sect. 3.3). It automatically excludes the training area from both the submitted and True source catalogs and produces a detection score for the rest of the image. Our reference score for this model is presented in Table. 5 under the name YOLO-CIANNA-ref, along with other scores from the top three originally participating teams on the same single 560 MHz - 1000h image. The Engage-SKA team used the ProFound package (Robotham et al., 2018), which is a collection of astronomical image processing and analysis tools that produce segmentation maps. The Shanghai team worked with a combination of three classical astronomical source-detector packages, namely AEGEAN (Hancock et al., 2018), DUCHAMP (Whiting, 2012), and SExtractor (Bertin & Arnouts, 1996), but ended up using only the last one to construct its catalog. Finally, the ICRAR team used a deep learning approach with the CLARAN method (Wu et al., 2019), with a region-based method similar to R-CNN (Girshick et al., 2013). We note the participation of another team that used a very shallow convolutional network structure for segmentation, achieving a much lower score with the ConvoSource method (Lukic et al., 2019). We also include another post-challenge score from Yu et al. (2022), which is a refinement of the JLRAT team that participated in the original challenge but with a lower submitted score, also with a deep learning-based method using a combination of a region-based approach with a feature pyramid network. Their new catalog is publicly available and is provided in the scorer format, allowing us to include their result on the same image in the table. Our reference score catalog is also made publicly available with the present paper so the full table can be reproduced. In addition to the challenge score, we also indicate the number of sources retrieved, the precision of the submitted catalog, and the average source characterization score.
The source catalog obtained with our YOLO-CIANNA detector improves the best-submitted score of the originally participating teams by +139% and the score of the only other published post-challenge score by +61%, which already significantly outperformed what was submitted during the challenge. We observe that the teams that competed in the challenge and obtained good scores are the ones that focused on catalog purity. This is likely to be due to their relatively low average source characterization score, which is discussed in Sect. 4.4. As the apparent flux should be an appropriate first-order proxy to represent the detection difficulty, we present the apparent flux distribution of our detections in comparison to the True catalog in Fig. 18. We also present the completeness and purity of our catalog for each flux bin. The completeness represents the proportion of total sources we successfully detected, and the purity represents the proportion of the proposed detection that are real sources. These quantities are computed based on the full True catalog and SDC1 matching criteria. Comparing this flux distribution with the one from our selected training sample (Fig. 8), we observe that our detector can identify sources outside its training flux range close to the noise limit. This is likely to correspond to sources for which the perceived apparent flux was increased by the local noise or blended faint sources that add their flux so the network detects them as a single brighter source (Sect. 3.4). These two cases will result in sources with a target apparent flux close to or below the detection limit to be detected. This out-of-range detection capability is clearly visible in the example fields from Fig. 20. We also see that the completeness is high in the intermediate apparent flux range but drops progressively at lower flux values. While it can be explained by sources getting under the detection noise limits, it can also come from sources too blended to be detected individually. In such a context, the prediction will be associated with the brightest target, removing mostly low apparent flux sources. These missed sources are also visible in the example fields from Fig. 20. On the other side of the flux range, our detector does not recover a few bright sources, which can be explained by the absence of similar examples in our training sample. While they are too few to influence the scoring by their absence, they can still have a strong effect by adding false positives, which is discussed in Sects. 3.9 and 4.1.
@@ 4.3 Alternative purity focused detection
By trying to optimize for the best possible SDC1 score, we are forced toward a predefined catalog purity. In fact, our per-source characterization score defines the ratio of true-to-false detection above which the detections add score points (Sect. 3.3). This relation is not linear, as objects that are difficult to detect will likely be difficult to characterize. This is why we optimize the score using the ordered detection in objectness, which reflects the network confidence (Sect. 3.9). If catalog purity is the main concern, we can adjust the objectness threshold selection to select more confident detections without retraining the detector. With this approach, we can evaluate the performance of a given trained detector in different confidence regimes. We define the YOLO-CIANNA-alt catalog, obtained by searching for objectness thresholds that enforce a 99% purity on the same network. The corresponding score is presented in Table. 5. This purity-focused catalog produces a score that is 13% lower than our reference result but still much higher than other submissions. We observe that the average source score also increases with this selection, which confirms that the predicted objectness captures information about both the detection and characterization difficulty of the sources.
We represent the apparent flux distribution of this purity focus catalog in Fig. 18. Comparing this distribution with our reference result shows that the detections at low and high apparent fluxes get removed first. While the link between low flux and difficulty is straightforward, it is less clear why bright objects get removed more than intermediate apparent flux regimes. Our interpretation would be that the small number of bright source examples in the training sample leads to poor characterization, resulting in lower detection confidence. By looking at the distribution of false positives over the full image, we observe a much more homogeneous distribution than with the YOLO-CIANNA-ref model. Therefore, the optimized set of objectness thresholds for the purity-focused model seems to be driven by regions of the image that are in the intermediate beam regime that is interpolated between our training regions.
As discussed in Sect. 2.2, having a confidence score that is mapped by a continuous function means that the network samples all its sensitivity regimes simultaneously. From a user perspective, a given trained model can be adapted to various applications with either completeness or purity requirements. We discuss how this principle could be used to design an alternative detection metric in Sect. 5.2.
@@ 4.4 Source characterization
Figure 19: 2D histograms of the different parameter errors defined by the scorer as a function of relevant difficulty proxies. The left column represents the 2D histograms. The right column represents a 1D histogram of the relative error projected over the same axis as the left column. To the right of this projection and using the same axis projection, we indicate the specific response of the score function for each parameter following Eq. 19 and Fig. 6. The red dashed line indicates the edges of the area for which the parameter score is saturated at its maximum value. The orange lines and arrow indicate the reference values that are represented by the training sample. The gray overlay represents the area of the distribution for sources that are smaller than the beam size.
@@
In addition to the global score, the scorer code can be configured to produce detailed results, including the list of sources that are considered a match, along with the predicted and target properties for each of them. In the following sections, we use these products to perform a detailed analysis of the characterization capabilities of our YOLO-CIANNA detector.
@@
A primordial aspect of our produced source catalog is its high source characterization performance with $\bar{s}=0.7703$. Before analyzing it in detail, it is important to highlight that parameter prediction accuracy is impossible to decorelate from the detection capability. The average parameter score ignores the fact that the characterization is not equally difficult for all sources. For example, we could choose to detect the 10 sources that are the easiest to characterize and obtain a very high characterization score. In contrast, a method that detects many sources, including some that are difficult to characterize, would obtain a lower average characterization score even if it is better at characterizing the specific 10 easiest sources. This effect is well illustrated by the increase in average source score for our alternative purity-focused results compared to our reference score result in Table 5. Overall, this average source score metric penalizes more efficient source detectors that properly detect more difficult sources. To overcome this limitation, it would be possible to order or bin the sources by difficulty before evaluating the characterization capability of different methods on each bin. However, it seems impossible to define a characterization or detection difficulty proxy that would be method agnostic. The best yet incomplete proxies we have are source brightness or signal-to-noise ratio (SNR) measurements. A better solution to compare method characterization capabilities would be to use subsets of sources that are detected by the different methods. In our opinion, this is a significant limitation of the analysis from both SDC1 and SDC2 summary papers.
@@
Our approach to overcome this difficulty is to use 2D histograms of the parameter errors from sources considered as matches by the scorer as a function of a difficulty proxy, which is presented in Fig. 19. This figure also comprises histograms of the sources regarding only the parameter errors, which are directly aligned with their respective score response functions to ease the comparison. The following sub-sections detail the prediction capability for each type of predicted parameter.
@@
We achieve a very high positioning score of $\bar{s}^{pos}=0.964$, which is, in fact, a requirement for reaching a high detection recall and precision regarding the scorer match criteria. For point sources to be considered a match, we must reach sub-pixel positioning accuracy (Sect. 3.3). This typically explains why we configured our detector with a strong focus on the position loss in Table 2. From this point of view, the 0.3 error threshold for the position seems quite permissive and is not very useful in comparing methods regarding their positioning accuracy.
@@
Regarding the flux prediction, the distribution mostly centers on the score function response, granting a good average flux score of $\bar{s}^{flux}=0.646$. As for the position, sources with a bad flux prediction are more likely to be rejected based on the scorer match condition (Sect. 3.3), imposing practical limits on the accessible flux dispersion. For faint sources, the detector overestimates the flux, which is typical of regression prediction when generalizing below the training range limit. Overall, the detector tends to predict the minimum flux value it was trained on for any fainter objects. Still, we note that the faintest sources are only detectable when the local noise increases their apparent flux. This can result in a selection effect that is likely to bias the network training to overestimate the underlying flux of faint sources. This can explain the spreading of the flux predictions above the minimum training value prediction. For the few bright sources, we see that our detector underestimates the predicted flux, which is induced by our maximum value at training time. We also note that the pixel cut limits combined with the dynamic rescale we applied to our input image could contribute to these flux prediction limits (Sect. 3.7.1). Still, trying other normalization cuts did not result in significant flux prediction improvements.
@@
Regarding $Bmaj$ and $Bmin$, the global predictions are convincing and centered on the score response function with $\bar{s}^{Bmaj}=0.627$ and $\bar{s}^{Bmin}=0.594$, but we observe the same type of out-limit generalization behaviors. Here, these effects are reinforced by the fact that the True values to be predicted are not convolved with the synthesized beam. In practice, the detector is expected to predict non-available information for most point sources that are actually detectable. Still, we observe that predictions for $Bmin$ are surprisingly not completely collapsed toward a single solution below the synthesized beam size. In practice, we have adjusted our clipping values to reflect this capability that was observed during our first testing, so it is now at the best limit we could reach. For $Bmaj$ specifically, we observe that the predictions are distributed over two independent prediction regimes that both follow the same type of error function. This separation is induced by the source distribution over different detection units based on their size. This creates a discontinuity as only detection units with a large size prior are properly constrained to predict large $Bmaj$ values, while it is the opposite for detection units with small size priors. We note that this effect is strongly mitigated by the addition of the random box association regularization in our association function (Appendix A.6 and Fig. 4) that forces detection units to generalize outside of their size range. By forcing more detection unit independence, the different size regimes get more separated in the figure.
@@
For the position angle, the average score is surprisingly high with $\bar{s}^{PA}=0.601$, considering that most detected sources are smaller than the beam resolution limit. From the distribution, we observe that the prediction is not too noisy and that small sources distribute evenly over the error range, while the prediction accuracy improves progressively with the increase in source size.

### v2

@@ 4 Results and Analysis
In this section, we present the results of our best model over the full field of the selected SDC1 image. We stress that there are small variations of results from retraining a network with the same training setup, with typical variations of the best achievable score around ±$\pm 0.5\%$. In practice, the detailed analysis of this specific training holds perfectly with any training we have done using the same setup. The results in this section are ordered following the post-processing pipeline from the raw predicted quantities up to the predicted source catalog using the challenge format. The validation loss curve as a function of the training iteration is presented in Fig. 13. This figure also indicates the optimized score at a regular interval of 200 iterations. The following results are presented for iteration 3000, which achieved the best score for this specific training.
@@ 4.1 Patch-divided prediction results
From the raw model prediction, after applying the first objectness filtering and the in-patch NMS, we can obtain a list of all the detected sources per patch. We represent the respective number of sources for all patches over the whole image field in Fig. 14. We note that some sources contribute to multiple bins in this representation because the inter-patch NMS has not yet been applied. We see that the detector sensitivity follows the primary beam imprints over the full image. We do not observe a clear overdensity of detected sources in the training area region, which tends to indicate the absence of overtraining.
@@
Figure 15: Distribution of the number of detections in a single cell (nonempty only) after objectness filtering and in-patch NMS.
@@
To analyze in detail the behavior of the individual detection units from our detector output (Sect. 2.5), we represent a histogram of the number of found-useful detections per grid cell for all patches in Fig. 15. Here, empty grid cells are excluded, but they represent the most common case. We observe that cases with six or more detections are very few, four or five detections are not common but significant, and most nonempty grid cells contain less than three detections. Further analysis indicates that only around $26\%$ of the nonempty grid cells have more than one detection. This proportion drops to $4.5\%$ for more than two detections. In practice, the in-patch NMS process removes $78\%$ of the raw good detections. This indicates that many detection units are in use simultaneously and capable of detecting real sources, but they often end up detecting the same sources. Still, we observed that lowering the number of detection units lowers the score, indicating that they are useful for some specific contexts or that having more detection units better distributes the information. To further verify this idea, we represent in Fig. 16 a histogram of the number of predictions per detection unit for the whole image after the in-patch NMS. The distribution over the different detection units from the small-size regime is relatively homogeneous, confirming that all detection units are useful. We also observe that the relative proportions of each size regime are well preserved regarding the theoretical target distribution from Fig. 12, which confirms that our association process properly distributes the information over the available size regimes.
To better understand the representation regime of individual detection units, we can look at their respective distribution of predicted sizes and aspect ratios in Fig. 17. We observe that some boxes are more specialized regarding aspect ratios, which is natural for the intermediate size regime due to the two different ratios of size priors. The behavior of the small-size regime detection units is more surprising, as they all appear to prefer a positive aspect ratio when predicting small boxes and a negative aspect ratio when predicting large boxes. The cause of this behavior remains unidentified for the moment, but it is reproducible in all of the backbone and hyperparameter configurations we tried. Finally, the detection unit from the large-size regime and usually one or two from the small-size regime also predict sizes representative of the intermediate regime. Due to the difference in aspect ratio, only one of the intermediate-size regime units is likely to be used in a given grid cell. When two or more sources of intermediate size have to be detected simultaneously, a small or large box with a square prior might be better than using the wrong aspect ratio from the appropriate size regime.
Figure 17: Histograms of the predicted heights, widths, and aspect ratios for every detection unit. The object count and size axes are in log scale. The aspect ratio is computed as $(w-h)/\max{(w,h)}$. The number of predictions and the proportion of strictly positive and negative aspect ratios are also given for each detection unit. The red vertical lines indicate the width and height of the size priors.
@@ 4.2 Full catalog scoring
After the inter-patch NMS filtering and conversion of all the predicted source parameters, we obtain a source catalog for the full image that can be put through the SDC1 scorer (Sect. 3.3). It automatically excludes the training area and produces a detection score for the rest of the image. Our score is presented in Table 5 under the name MINERVA (YOLO-CIANNA), along with scores from other methods. In addition to the challenge score, we also indicate the number of sources retrieved, the precision of the submitted catalog, and the average source characterization score. For a fair comparison, we recomputed the scores from the available detection catalogs for each method using the same scoring code. The top three results from the original challenge leaderboard on the same single 560 MHz - 1000h image are indicated for reference. The Engage-SKA team used the ProFound package (Robotham et al., 2018), a collection of astronomical image processing and analysis tools that produce segmentation maps. The Shanghai team worked with a combination of three classical astronomical source-detector packages, namely AEGEAN (Hancock et al., 2018), DUCHAMP (Whiting, 2012), and SExtractor (Bertin & Arnouts, 1996), but ended up using only the last one to construct its catalog. Finally, the ICRAR team used the deep learning CLARAN method (Wu et al., 2019), with a region-based method similar to R-CNN (Girshick et al., 2013). We note the participation of another team (not in our table) that used a very shallow convolutional network structure for segmentation, achieving a much lower score with the ConvoSource method (Lukic et al., 2019). We also include another post-challenge score from Yu et al. (2022), which is a refinement of the JLRAT team that participated in the original challenge but with a lower submitted score, also with a deep learning-based method using a combination of a region-based approach with a feature pyramid network. Their new catalog is publicly available and is provided in the scorer format, allowing us to include their result on the same image in the table. Our SDC1 source catalog is also made publicly available (Cornu, 2024c) [^3] Our reference SDC1 detection source catalog along with other interesting catalogs described in the present paper have been archived at 10.5281/zenodo.13141772 in the scorer format with the present paper so the full score table can be reproduced.
The source catalog obtained with our YOLO-CIANNA detector improves the best-submitted score of the originally participating teams by +139% and the score of the only other published post-challenge score by +61%, which already outperformed what was submitted during the challenge. We observe that most teams focused on catalog purity to achieve high scores. This is likely to be due to their relatively low average source characterization score, which is discussed in Sect. 4.4. As the apparent flux should be an appropriate first-order proxy to represent the detection difficulty, we present the apparent flux distribution of our detection list in comparison to the True catalog in Fig. 18. We also present the corresponding completeness and purity for each flux bin. The completeness represents the proportion of total sources we successfully detected, and the purity represents the proportion of the proposed detection that are real sources. These quantities are computed based on the full True catalog and SDC1 matching criteria. Comparing this flux distribution with the one from our selected training sample (Fig. 8), we observe that our detector can identify sources outside its training flux range close to the noise limit. This is likely to correspond to sources for which the perceived apparent flux was increased by the local noise or blended faint sources that add their flux, leading the network to detect them as a single brighter source (Sect. 3.4). These two cases will result in sources with a target apparent flux close to or below the detection limit to be detected. This out-of-range detection capability is also visible in the example fields from Fig. 20. The completeness is high in the intermediate apparent flux range but drops progressively at lower flux values. While it can be explained by sources getting under the detection noise limits, it can also come from sources that are too blended to be detected individually. In such a context, the prediction will be associated with the brightest target, removing mostly low apparent flux sources. This is also visible in the example fields from Fig. 20. On the opposite side of the flux range, our detector misses a few bright sources, which can be explained by the absence of similar examples in our training sample. While they are too few to influence the scoring by their absence, they can still have a strong effect by adding false positives, which is discussed in Sects. 3.9 and 4.1.
Table 5: . SDC1 scores and related properties for source catalogs from different teams and methods.
Team (method) $M_{s}$ (Score) $N_{\rm det}$ $N_{\textrm{match}}$ $N_{\rm false}$ $N_{\rm bad}\in N_{\rm false}$ Purity $\bar{s}$
@@
MINERVA (YOLO-CIANNA) 480450 724480 680000 44480 16839 93.86% 0.7719
$\hookrightarrow$ purity-focus thresholds 418434 541542 536412 5130 2506 99.06% 0.7896
@@
Score-focus selection
Purity-focus selection
Figure 18: Histograms of the sources as a function of their flux using logarithmic bins for the underlying True catalog and the predicted sources in the testing area. The true source flux is used for matches, while the predicted flux is used for false positives. The bottom part of each frame represents the purity and completeness of each bin of the histograms (defined in Sect. 4.2). The top frames use the apparent flux after multiplication with the primary beam, while the bottom frames use the integrated flux. The columns represent the two possible post-process objectness-thresholds search objectives over the same trained YOLO-CIANNA model.
@@ 4.3 Alternative purity focused detection
By trying to optimize for the best possible SDC1 score, we are forced toward a predefined catalog purity. Consequently, our per-source characterization score defines the ratio of true-to-false detections above which new detections add score points (Sect. 3.3). This relation is nonlinear, as objects that are difficult to detect will likely be challenging to characterize. This is why we optimize the score using the ordered detection in objectness, which reflects the detector confidence (Sect. 3.9). If catalog purity is the main concern, we can adjust the objectness threshold selection to preserve only more confident detections using the exact same model (no retraining is necessary). With this approach, we can evaluate the performance of a given trained detector in different confidence regimes. To illustrate this capability, we present the result obtained with the same trained YOLO-CIANNA model but with a post-process search for objectness thresholds that enforce a 99% purity. The score for this purity-focus catalog is presented in Table 5 as a subproduct of our model. As expected, keeping only the most confident sources reduces the score by around 13% compared to our score-optimized catalog, as it removes sources that were contributing a positive score. Still, this score remains much higher than the other methods with a similar purity. We observe that the average source score also increases with this selection, which confirms that the predicted objectness captures information about both the detection and characterization difficulty of the sources.
We represent the apparent flux distribution of this purity focus catalog in Fig. 18. Comparing this distribution with our default result shows that the detections at low and high apparent fluxes get removed first. While the link between low flux and difficulty is straightforward, it is less clear why bright objects get removed more than intermediate apparent flux regimes. It is likely due to the small number of bright source examples in the training sample that leads to poor characterization, resulting in lower detection confidence. It can also be related to the fact that extended sources often exhibit bright-compact regions over much fainter extended structures, resulting in target bounding boxes that do not correlate well with the detectable part of the objects. This naturally prevents the detector from predicting the appropriate box sizes and lowers the geometrical confidence reflected in the objectness of extended sources.
As discussed in Sect. 2.2, having a predicted confidence score mapped by a continuous function implies that the trained detector can sample all its sensitivity regimes simultaneously at prediction time. By optimizing for the best SDC1 score or for a given purity, we only select specific sensitivity thresholds. From a user perspective, a given trained model can be adapted to various applications with either completeness or purity requirements. We discuss how this principle could be used to design an alternative detection metric in Sect. 5.2.
@@ 4.4 Source characterization
Figure 19: Two-dimensional histograms of the different parameter errors defined by the scorer as a function of relevant difficulty proxies. The left column represents the two-dimensional histograms. The right column represents a 1D histogram of the relative error projected over the same axis as the left column. To the right of this projection and using the same axis projection, we indicate the specific response of the score function for each parameter following Eq. 19 and Fig. 6. The red dashed line indicates the edges of the area for which the parameter score is saturated at its maximum value. The orange lines and arrow indicate the value range available in the training sample. The gray overlay represents the size regime for which the sources are smaller than the beam size.
@@
In addition to the global score, the scorer code can produce detailed results, which include a list of matches along with the predicted and target properties for each. In the following sections, we use these products to perform a detailed analysis of the characterization capabilities of our YOLO-CIANNA detector.
@@
A primordial aspect of our produced source catalog is its high source characterization performance with $\bar{s}=0.7703$. Before analyzing it in detail, it is important to highlight that parameter prediction accuracy is impossible to decorelate from the detection capability in the SDC1 scorer. The average parameter score ignores the fact that the characterization is not equally difficult for all sources. For example, we could choose to detect only the ten sources that are the easiest to characterize and obtain a very high characterization score. In contrast, a method that detects many sources, including some that are difficult to characterize, would get a lower average characterization score even if it is better at characterizing the specific ten easiest sources. This effect is well illustrated by the increase in average source score for our alternative purity-focused result compared to our default score result in Table 5. Overall, this average source score metric penalizes more efficient detectors that properly identify more difficult sources. To overcome this limitation, it would be possible to order or bin the sources by difficulty before evaluating the characterization capability of different methods on each bin. However, defining a characterization or detection difficulty proxy that would be method agnostic is likely impossible. The best proxies we have are source brightness or signal-to-noise ratio (S/N) measurements. A better solution to compare method characterization capabilities would be to use subsets of sources detected by the different methods. In our opinion, this is a significant limitation of the analysis from both SDC1 and SDC2 summary papers.
@@
Our approach to overcome this difficulty is to use 2D histograms of the parameter errors from sources matched by the scorer as a function of a difficulty proxy, presented in Fig. 19. This figure also comprises histograms of the sources regarding only the parameter errors, which are directly aligned with their respective score response functions to ease the comparison. The following subsections detail the prediction capability for each type of predicted parameter.
@@
We achieve a very high positioning score of $\bar{s}^{\textrm{pos}}=0.964$, which is, in fact, a requirement for reaching a high detection recall and precision regarding the scorer match criteria. For point sources to be considered a match, we must reach subpixel positioning accuracy (Sect. 3.3). This typically explains why we configured our detector with a strong focus on the position loss in Table 2. From this point of view, the 0.3 error threshold for the position seems relatively permissive and is not very useful in comparing methods regarding their positioning accuracy.
@@
For the flux prediction, the distribution mostly centers on the score function response, granting a good average flux score of $\bar{s}^{\textrm{flux}}=0.646$. As for the position, sources with a bad flux prediction are more likely to be rejected based on the scorer match condition (Sect. 3.3), imposing practical limits on the accessible flux dispersion. For faint sources, the detector overestimates the flux, which is typical of a prediction over a noise-limited image. Overall, the detector tends to predict the minimum flux value it was trained on for fainter objects. Still, the faintest sources are only detectable when the local noise increases their apparent flux. This can result in a selection effect bias that causes the network to overestimate the underlying flux of faint sources. This can explain the spreading of the flux predictions above the minimum training value prediction. For the few bright sources, our detector underestimates the predicted flux, which results from the combination of several factors: i) our maximum value at training time will push the network to underestimate all fluxes above the limit, ii) high flux sources are represented mainly by a rare class of objects (Sect. 3.4) that are only detectable at this flux regimes meaning that they are unconstrained, and iii) our hyperbolic tangent transform applied in the input renormalization might slightly reduce the accessible information for high flux sources (Sect. 3.7.1).
@@
Regarding Bmaj and Bmin, the global predictions are convincing and centered on the score response function with $\bar{s}^{\textrm{Bmaj}}=0.627$ and $\bar{s}^{\textrm{Bmin}}=0.594$. We also observe the same type of out-limit generalization behaviors. Here, these effects are reinforced by the fact that the True values to be predicted are not convolved with the synthesized beam. In practice, the detector is expected to predict nonavailable information for most detectable point sources. Still, we observe that predictions for Bmin are surprisingly not collapsed toward a single solution below the synthesized beam size. This indicates that the network can extract residual information about the source size under the resolution limit or reconstruct this information from correlated parameters. For Bmaj specifically, we observe that the predictions are distributed over two independent prediction regimes that both follow the same type of error function. This separation is induced by the source distribution over different detection units based on their size. This creates a discontinuity as only detection units with a large size prior are properly constrained to predict large Bmaj values, while small detection units with small size priors are constrained to predict small Bmaj values. This effect is already strongly mitigated by the addition of the random box association regularization in our association function (Appendix A.6 and Fig. 5) that forces detection units to generalize outside of their size range. When forcing more detection unit independence, this separation between different size regimes gets even more striking in the figure.
@@
For the position angle, the average score is surprisingly high with $\bar{s}^{\textrm{PA}}=0.601$, considering that most detected sources are smaller than the beam resolution limit. From the distribution, we observe that the prediction is not too noisy and that small sources distribute evenly over the error range, while the prediction accuracy improves progressively with the increase in source size.